What is the equation of a parabola with vertex (0, 0) and directrix x = -3.
Answer by DSMLMD(16)

Vertex at (0, 0) and directrix x = -3


Vertex
(0,0)
(a,b)


directrix formula:
x = -p
x = -(-3)
x = 3


because the directrix located at x-coordinate, the parabola equation is:


(y-b)^2 = 4p(x-a)
(y-0)^2 = 4(3)(x-0)
y^2 = 12x